ATK Tähtitieteessä 11.04.2001

1. IDL-proseduurit Pääohjelma Aliohjelmat

PRO aliohjelma,arg1,arg2,...,argN,keyword1=keyword1,keyword2=keyword2,... Funktiot

FUNCTION funktion_nimi,arg1,arg2,...,keyword1=keyword1,keyword2=keyword2,...

2. Komentorakenteet For-silmukka If-rakenne While-rakenne

3. PS-tiedoston luonti

4. 2D-plotit

5. Harjoitustöiden jako

ATK Tähtitieteessä 11.04.2001

1. IDL-proseduurit Pääohjelma Aliohjelmat

PRO aliohjelma,arg1,arg2,...,argN,keyword1=keyword1,keyword2=keyword2,... Funktiot

FUNCTION funktion_nimi,arg1,arg2,...,keyword1=keyword1,keyword2=keyword2,...

2. Komentorakenteet For-silmukka If-rakenne While-rakenne

3. PS-tiedoston luonti

4. 2D-plotit

5. Harjoitustöiden jako